# Intuitionistic NF set theory

### v5

@@ 12. Onto and one-to-one for maps between finite sets
That completes the induction step. That completes the proof of the lemma.
13. The initial segments of ${\mathbb{F}}$
@@
Proof. By the definitions of ${\mathbb{J}}$ and $\bar{\mathbb{J}}$, and the fact that for $x\in{\mathbb{F}}$ we have
@@
Proof. By induction on $m$. The formulas to be proved, namely
$\forall m\,(m\in{\mathbb{F}}\ {\rightarrow}\ {\mathbb{J}}(m)\in{\sf FINITE}$
@@
Base case, $m={\sf zero}$. Then ${\mathbb{J}}({\sf zero})=\Lambda$, by Lemma 5.33. By Lemma 3.6, $\Lambda\in{\sf FINITE}$. That completes the base case for ${\mathbb{J}}$. For $\bar{J}$, we have $x\leq{\sf zero}{\ \leftrightarrow\ }x={\sf zero}$, so $\bar{J}({\sf zero})=\{{\sf zero}\}$, which is finite by Lemma 3.7. That completes the base case.
Induction step. Suppose $m\in{\mathbb{F}}$ and $m^{+}$ is inhabited. By induction hypothesis, ${\mathbb{J}}(m)$ and $\bar{\mathbb{J}}(m)$ are finite. By Lemma 13.2, ${\mathbb{J}}(m^{+})={\mathbb{J}}(m)\cup\{m\}$, so by Lemma 3.7, $J(m^{+})\in{\sf FINITE}$. Similarly for $\bar{\mathbb{J}}(m)$. That completes the induction step. That completes the proof of the lemma.
@@
Suppose $m\in{\mathbb{F}}$. Then $Nc({\mathbb{J}}(m))={\mathbb{T}}^{2}m.$
Proof. The formula of the lemma is stratified, giving $m$ index 0, since then ${\mathbb{T}}^{2}m$ gets index 2, while ${\mathbb{J}}(m)$ gets index 1 and $Nc({\mathbb{J}}(m))$ gets index 2, so the two sides of the equation both get index 2. Therefore the lemma may be proved by induction.
Base case: ${\mathbb{J}}({\sf zero})=\Lambda$, by Lemma 5.28. We have $Nc(\Lambda)={\sf zero}$, by Lemma 4.11 and the definition of ${\sf zero}$. By Lemma 10.9, we have ${\mathbb{T}}^{2}{\sf zero}={\sf zero}$. That completes the base case.
@@
$Nc({\mathbb{J}}(m))={\mathbb{T}}^{2}m$   by the induction hypothesis
@@
${\mathbb{J}}(m^{+})\in Nc({\mathbb{J}}(m^{+}))$   by Lemma 4.11
${\mathbb{J}}(m^{+})\in{\mathbb{T}}^{2}(m^{+})\cap Nc({\mathbb{J}}(m^{+}))$   by definition of intersection
$Nc({\mathbb{J}}(m^{+}))={\mathbb{T}}^{2}(m^{+})$   by Lemma 4.7
That completes the induction step. That completes the proof of the lemma.
14. Rosser’s Counting Axiom
@@
Since $2^{{\mathbb{T}}m}$ is always defined for $m\in{\mathbb{F}}$, the counting axiom implies that $2^{m}$ is always defined for $m\in{\mathbb{F}}$. In particular then the set of iterated powers of $2$ starting from ${\sf zero}$ is an infinite set. That is the conclusion of Specker’s proof (but without assuming the counting axiom). The point here is that the counting axiom eliminates the need to constructivize Specker’s proof: if we assume it, there remain only surmountable difficulties to interpreting HA in INF. But the counting axiom is stronger than NF [7], so this observation does not help with the problem of finiteness in INF.

### v6

@@ 12 Onto and one-to-one for maps between finite sets
That completes the induction step. ∎
13 The initial segments of $\mathbb{F}$}
@@
Proof.
By the definitions of $\mathbb{J}$} and $\bar{\mathbb{J}}$, and the fact that for $x\in{\mathbb{F}}$ we have
@@
Proof.
By induction on $m$. The formulas to be proved, namely
$\forall m\,(m\in{\mathbb{F}}\ {\rightarrow}\ {\mathbb{J}}(m)\in{\sf FINITE})$
@@
Base case, $m={\sf zero}$. Then ${\mathbb{J}}({\sf zero})=\emptyset$, by Lemma 5.34. By Lemma 3.6, $\emptyset\in{\sf FINITE}$. That completes the base case for $\mathbb{J}$}. For $\bar{J}$, we have $x\leq{\sf zero}{\ \leftrightarrow\ }x={\sf zero}$, so $\bar{J}({\sf zero})=\{{\sf zero}\}$, which is finite by Lemma 3.7. That completes the base case.
Induction step. Suppose $m\in{\mathbb{F}}$ and $m^{+}$ is inhabited. By induction hypothesis, ${\mathbb{J}}(m)$ and $\bar{\mathbb{J}}(m)$ are finite. By Lemma 13.2, ${\mathbb{J}}(m^{+})={\mathbb{J}}(m)\cup\{m\}$, so by Lemma 3.7, $J(m^{+})\in{\sf FINITE}$. Similarly for $\bar{\mathbb{J}}(m)$. That completes the induction step. ∎
@@
Suppose $m\in{\mathbb{F}}$. Then $|{\mathbb{J}}(m|)={\mathbb{T}}^{2}m$.
Proof.
The formula of the lemma is stratified, giving $m$ index 0, since then ${\mathbb{T}}^{2}m$ gets index 2, while ${\mathbb{J}}(m)$ gets index 1 and $|{\mathbb{J}}(m|)$ gets index 2, so the two sides of the equation both get index 2. Therefore the lemma may be proved by induction.
Base case: ${\mathbb{J}}({\sf zero})=\emptyset$, by Lemma 5.29. We have $|\emptyset|={\sf zero}$, by Lemma 4.11 and the definition of $\sf zero$}. By Lemma 10.9, we have ${\mathbb{T}}^{2}{\sf zero}={\sf zero}$. That completes the base case.
@@
$|{\mathbb{J}}(m|)={\mathbb{T}}^{2}m$   by the induction hypothesis
@@
${\mathbb{J}}(m^{+})\in|{\mathbb{J}}(m^{+}|)$   by Lemma 4.11
${\mathbb{J}}(m^{+})\in{\mathbb{T}}^{2}(m^{+})\cap|{\mathbb{J}}(m^{+}|)$   by definition of intersection
$|{\mathbb{J}}(m^{+}|)={\mathbb{T}}^{2}(m^{+})$   by Lemma 4.7
That completes the induction step. ∎
14 Rosser’s Counting Axiom
@@
Since $2^{{\mathbb{T}}m}$ is always defined for $m\in{\mathbb{F}}$, the counting axiom implies that $2^{m}$ is always defined for $m\in{\mathbb{F}}$. In particular then the set of iterated powers of $2$ starting from $\sf zero$} is an infinite set. That is the conclusion of Specker’s proof (but without assuming the counting axiom). The point here is that the counting axiom eliminates the need to constructivize Specker’s proof: if we assume it, there remain only surmountable difficulties to interpreting HA in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. But the counting axiom is stronger than NF [13], so this observation does not help with the problem of finiteness in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
15 Infinity in intuitionistic NF
We use Dedekind’s definition, that a set is infinite if it is similar to a proper subset. The “axiom of infinity” says there is an infinite set. Before going further, we remind the reader that with intuitionistic logic, “not finite” does not imply “infinite”. There are two obvious candidates for infinite sets: $\mathbb{V}$} and $\mathbb{F}$}. Specker showed that, with classical logic, $\mathbb{V}$} is not finite; we will discuss that proof below.
If $\mathbb{F}$} is finite, then by Lemma 5.35, there is a maximal finite cardinal $\bf m$}. Then by Lemma 4.7, $\bf m$} has a member $U$, and by Lemma 4.21, $U$ is finite. If we could find some $c\not\in U$, then $\bf m$$^{+}$ would be inhabited and hence in $\mathbb{F}$}, contradicting the maximality of $\bf m$}. Therefore $\forall x\,\neg\neg\,(x\in U)$; that is, $\mathbb{V}$} is the double complement of $U$. However unlikely this may seem, nobody has yet been able to find anything contradictory about it, without using classical logic. The following lemma states this remarkable result, so we can cite it below.
Lemma 15.1.
Suppose $\bf m$} is a maximal element of $\mathbb{F}$}, and $U\in{\bf m}$. Then $\forall x\,\neg\neg\,(x\in U)$.
Lemma 15.2.
Let $\bf m$} is a maximal element of $\mathbb{F}$} and $n\in{\mathbb{F}}$. Then ${\mathbb{T}}{\bf m}<n$ implies $2^{n}=\emptyset$.
Proof.
Suppose ${\mathbb{T}}{\bf m}<n$ and $2^{n}$ is inhabited; we must derive a contradiction.
$\mathcal{P}_{1}(u)\in n$   for some $u$, by definition of exponentiation
$u\in|u|$   by Lemma 4.11
$\mathcal{P}_{1}(u)\in{\mathbb{T}}(|u|)$   by definition of $\mathbb{T}$}
${\mathbb{T}}n={\mathbb{T}}(|u|)$   by Lemma 4.24
${\mathbb{T}}{\bf m}<{\mathbb{T}}(|u|)$   since ${\mathbb{T}}{\bf m}<n$
${\bf m}<|u|$   by Lemma 10.20
But that contradicts the maximality of $\bf m$}. ∎
Lemma 15.3.
If $\mathbb{V}$} is infinite then $\mathbb{F}$} is not finite.
Remark. Note that Specker’s proof shows $\mathbb{V}$} is not finite, but not that $\mathbb{V}$} is infinite, which is stronger.
Proof.
Suppose $\mathbb{V}$} is infinite and $\mathbb{F}$} is finite, with maximal integer $\bf m$} and $U\in{\bf m}$ and $f:{\mathbb{V}}\to{\mathbb{V}}$ with $c$ not in the range of $f$. Then
$\forall x\,\neg\neg\,x\in U$   since $U\in{\bf m}$
$\forall x\,(x\in U\ {\rightarrow}\ \neg\neg\,(f(x)\in U))$   by the previous line
$\neg\neg\forall x\,(x\in U\ {\rightarrow}\ f(x)\in U)$   by Lemma 3.28
$\neg\neg\,(f:U\to U)$   by definition of $f:U\to U$
$\neg\neg\,(c\in U)$   since $\forall x\,\neg\neg\,x\in U$
That implies that $U$ is not not infinite. But since $U$ is finite, it is not infinite, by Theorem 3.24. ∎
Lemma 15.4.
With classical logic, if $\mathbb{V}$} is not finite then $\mathbb{F}$} is not finite.
Proof.
Suppose $\mathbb{V}$} is not finite and $\mathbb{F}$} is finite. Let $\bf m$} be the maximal integer and $U\in{\bf m}$. Then $U$ is finite and $\forall x\,\neg\neg\,(x\in U)$. Then by classical logic, ${\mathbb{V}}=U$, contradiction, since $U$ is finite and $\mathbb{V}$} is not. ∎
But constructively, the situation is more complicated: we can prove $\mathbb{V}$} is not finite, but it is an open problem whether $\mathbb{F}$} is finite or not.
To prove $\mathbb{F}$} is infinite, we would hope to prove that successor maps $\mathbb{F}$} into $\mathbb{F}$}, so it is of some interest whether that follows from the apparently weaker proposition that $\mathbb{F}$} is not finite. We cannot answer that question: it is an open problem whether
${\mathbb{F}}\in{\sf FINITE}\ {\rightarrow}\ \forall x\in{\mathbb{F}}\,(x^{+}\in{\mathbb{F}}).$
In other words, as far as we know, it might be that $\forall\,U\in{\sf FINITE}\,({\mathbb{V}}-U\neq\emptyset)$, but nevertheless we cannot prove $\forall\,U\in{\sf FINITE}\,\exists x\,(x\in{\mathbb{V}}-U)$. The former is equivalent to successor being nonempty on $\mathbb{F}$}, the latter to successor being inhabited on $\mathbb{F}$}. We cannot shift the double negation left through $\neg\neg$. (We shall see below that $\sf FINITE$} is not finite, so Lemma 3.28 is no use here.)
Nevertheless, if we did somehow prove that $\mathbb{F}$} is finite, we could prove that Heyting’s arithmetic HA is interpretable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. Here is how we would do that:
Recall that $\mathbb{F}$} is the least set containing $\sf zero$} and closed under inhabited successor. Now define $\mathbb{H}$} to be the least set containing $\sf zero$} and closed under nonempty successor. Then we can prove things using $\mathbb{H}$}-induction, in which at the induction step one is allowed to assume $x^{+}\neq\emptyset$, instead of the usual $\exists u\,(u\in x^{+})$. Assume that $\mathbb{F}$} is not finite. We do not give all the details, but here is a sketch: First we prove $\mathbb{F}$$\subseteq{\mathbb{H}}$, by $\mathbb{F}$}-induction. Then by $\mathbb{H}$}-induction, we prove $\forall x\in{\mathbb{H}}\,(\neg\neg\,x\in{\mathbb{F}})$, then $\emptyset\not\in{\mathbb{H}}$; then that $\mathbb{H}$} is closed under successor and has decidable equality, and that successor is one to one on $\mathbb{H}$}. Then we could use $\mathbb{H}$} as the interpretation of the variables of HA. But that would still not prove that $\mathbb{F}$} is closed under successor!
In the sea of open problems, there is an island: the theorem of Specker that $\mathbb{V}$} is not finite. This theorem, proved classically in [17], is widely acknowledged as constructively correct, for reasons I will now explain. Let $P$ be any stratified formula and let $X_{P}=\{x\in\{\emptyset\}:P\}$. Then $X_{P}$ is $\sf zero$} or $\emptyset$ according as $P$ or $\neg P$. If $\mathbb{V}$} is finite then $\mathbb{V}$} has decidable equality, so by deciding whether $X_{P}=\emptyset$ or not, we decide P¬PP\ $\lor\ \neg$ P. Then, folklore has it, Specker’s proof of infinity uses classical logic only for stratified formulas, so it will go through under the assumption that $\mathbb{V}$} is finite, and produce a contradiction.
While this metamathematical argument is appealing, it still requires checking the details of Specker’s proof to ensure that classical logic is used only for stratified formulas. I studied Specker’s proof, trying to make it constructive, and using Lean to check my proofs. Assume there is a maximal integer $\bf m$}. Then $\bf m$} has a member $U$, which is “unenlargeable”, as discussed above. I thought that perhaps $U$ could be made to play the role that $\mathbb{V}$} plays in Specker’s proof. That plan did not succeed, unless we assume $\mathbb{V}$} is finite, in which case Specker’s proof does provide a Lean-checkable proof that $\mathbb{V}$} is not finite. I chose not to present it here. [^12] It is not very short; the details are in no doubt; it leads to an even lengthier discussion of the problem of infinity, but not to a solution of that problem.
Rosser, in an appendix to [16] (but not the first edition [15]), gave another proof that $\mathbb{V}$} is not finite, in which Specker’s ideas are recognizable. Rosser proves $\mathbb{V}$} is not finite and then immediately concludes that $\mathbb{F}$} is not finite, since classically $m\in{\mathbb{F}}$ and $\in m, UU$ is finite so ${\mathbb{V}}-U$ is inhabited, so $m^{+}$ is inhabited. The proof that $\mathbb{V}$} is not finite might well be constructive. I did not check it in Lean, since I already checked Specker’s proof in Lean.
Once we know that $\mathbb{V}$} is not finite, we can try to prove other sets are not finite. For example, $\sf FINITE$} is not finite, as we shall prove soon.
Lemma 15.5.
$\forall x\,(x\in{\sf FINITE}\ {\rightarrow}\ x\in\mathcal{P}_{1}({\mathbb{V}})\ \lor\ x\not\in\mathcal{P}_{1}({\mathbb{V}}))$.
Proof.
A set $x$ is a singleton if and only if $|x|={\sf one}$. That is,
$\forall x\,(x\in\mathcal{P}_{1}({\mathbb{V}}){\ \leftrightarrow\ }|x|={\sf one})$   by the definitions of $\mathcal{P}_{1}$ and $\sf one$}
Since equality on $\mathbb{F}$} is decidable, it is decidable whether a finite set is a singleton or not. Therefore
$\forall x\,(x\in{\sf FINITE}\ {\rightarrow}\ x\in\mathcal{P}_{1}({\mathbb{V}})\ \lor\ x\not\in\mathcal{P}_{1}({\mathbb{V}}))$
Lemma 15.6.
$\sf FINITE$} is not finite.
Remark. This depends on the fact the $\mathbb{V}$} is not finite, which we do not list as a hypothesis, since it is a theorem, even if the proof has not been presented here.
Proof.
Assume $\sf FINITE$} is finite. We must derive a contradiction. We have
$\mathcal{P}_{1}({\mathbb{V}})\subseteq{\sf FINITE}$   by Lemma 3.9
$\mathcal{P}_{1}({\mathbb{V}})\in\mathcal{P}_{s}({\sf FINITE})$   by Lemma 15.5
$\mathcal{P}_{1}({\mathbb{V}})\in{\sf FINITE}$   by Lemma 3.19, since ${\sf FINITE}\in{\sf FINITE}$
${\mathbb{V}}\in{\sf FINITE}$   by Lemma 3.10
16 Conclusions
This paper lays the foundations for future studies of intuitionistic NF set theory $\mathit{i}\mkern 0.4mu\mathrm{NF}$, by providing coherent definitions for the basic concepts, including order, exponentiation, addition, finite sets, and $\mathbb{T}$}. The concept of separability plays an important role in order and power set, and hence in exponentiation as well. The theory presented here—if supplemented by a proof that the set of integers is not finite—would serve well as a basis for formalizing constructive mathematics in the style of Bishop. These basic theorems will surely be both useful and necessary for deeper investigations of the metamathematical properties of $\mathit{i}\mkern 0.4mu\mathrm{NF}$. That subject has yet to begin, as at present we cannot even show that the law of the excluded middle is not provable in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.